OpenStudy (anonymous):

2/3 of all your books are fiction. 1/6 of those are historical fiction. What fraction of all your books are historical fiction?

OpenStudy (ahsome):

Multiply the fractions together

OpenStudy (ahsome):

\[\frac{2}{3}\times \frac{1}{6}\]

OpenStudy (ahsome):

\[\frac{2}{18}\]\[\frac{1}{9}\]